(2x+1)(x+1)d^2y/dx^2+2xdy/dx-2y=(2x+1)^2
1
Expert's answer
2020-07-20T18:53:42-0400

"(2x+1)(x+1)y''(x)+2xy'(x)-2y=(2x+1)^2"  



Ordinary differential equation classification: second-order linear ordinart differential equation.


Step by step solution:

1-step: Substitute "-2= \\tfrac{d}{dx}\t(2x) - \\tfrac{d^2}{dx}((x+1)(2x+1)):"


"\\tfrac{d^2 y(x)}{dx^2}\t(x+1)(2x+1)+2x \\tfrac{dy(x)}{dx}+( \\tfrac{d}{dx}(2x)- \\tfrac{d^2}{dx^2}((x+1)(2x+1)))y(x) = (2x +1)^2"


2-step: Expand


"\tfrac{d^2 y(x)}{dx^2}(x+1)(2x+1)+2x\tfrac{dy(x)}{dx}+\tfrac{d}{dx}(2x)y(x)-\tfrac{d^2}{dx^2}((x+1)(2x+1))y(x)=(2x+1)^2"


3-step: Add and subtract "\tfrac{dy(x)}{dx}\tfrac{d}{dx}((x+1)(2x+1))" to the left hand sinde:


"\tfrac{dy(x)}{dx}\tfrac{d}{dx}((x+1)(2x+1))+\tfrac{d^2y(x)}{dx^2}(x+1)(2x+1)+2x\tfrac{dy(x)}{dx}+\tfrac{d}{dx}(2x)y(x)-\tfrac{dy(x)}{dx}\tfrac{d}{dx}((x+1)(2x+1))-\tfrac{d^2}{dx^2}((x+1)(2x+1))y(x)=(2x+1)^2"


4-step: Apply the reverse product rule "f\tfrac{dg}{dx}+g\tfrac{df}{dx}=\tfrac{d}{dx}(fg)" to the left-hand side:


"\tfrac{d}{dx}(\tfrac{dy(x)}{dx}(x+1)(2x+1))+\tfrac{d}{dx}(2xy(x))-\tfrac{d}{dx}((2x+2(x+1)+1)y(x))=(2x+1)^2"


5-step: Factor:


"\tfrac{d}{dx}(\tfrac{dy(x)}{dx}(x+1)(2x+1)+2xy(x)-(2x+2(x+1)+1)y(x))=(2x+1)^2"


6-step: Integrate both sides with respect to x:


"\int\tfrac{d}{dx}(\tfrac{dy(x)}{dx}(x+1)(2x+1)+2xy(x)-(2x+2(x+1)+1)y(x))dx=\int(2x+1)^2dx"


7-step: Evalute the integrals:


"\tfrac{dy(x)}{dx}(x+1)(2x+1)+2xy(x)-(2x+2(x+1)+1)y(x)=\tfrac{4x^3}{3}+2x^3+x+" C1

where C1 is an arbitrary constant.

8-step: Rewrite the equation:


"\tfrac{dy(x)}{dx}+\tfrac{-2(x+1)-1)y(x)}{(x+1)(2x+1)}=\tfrac{-\tfrac{4x^3}{3}-2x^3-x-C}{(x+1)(2x+1)}"


9-step: Let "\mu(x)=exp(\int\tfrac{-2(x+1)-1}{(x+1)(2x+1)}dx)= \tfrac{x+1}{(2x+1)^2}"

Multiply both sides by "\mu(x):"

"\tfrac{(x+1)\tfrac{dy(x)}{dx}}{(2x+1)^2}+\tfrac{(-2(x+1)-1)y(x)}{(2x+1)^3}=-\tfrac{-\tfrac{4x^3}{3}-2x^3-x-C}{(2x+1)^3}"


10-step: Substitute "\tfrac{-2(x+1)-1}{(2x+1)^3}=\tfrac{d}{dx}(\tfrac{x+1}{(2x+1)^2}):"


"\tfrac{(x+1)\tfrac{dy(x)}{dx}}{(2x+1)^2}+\tfrac{d}{dx}(\tfrac{x+1}{(2x+1)^2})y(x)=-\tfrac{-\tfrac{4x^3}{3}-2x^3-x-C}{(2x+1)^3}"


11-step: Apply the reverse product rule "f\tfrac{dg}{dx}+g\tfrac{df}{dx}=\tfrac{d}{dx}(fg)" to the ledt-hand side:


"\tfrac{d}{dx}( \tfrac{(x+1)y(x)}{(2x+1)^2}= -\tfrac{-\tfrac{4x^3}{3}-2x^3-x-C}{(2x+1)^3}"


12-step: Integrate both sides with respect to x:


"\int\tfrac{d}{dx}(\tfrac{(x+1)y(x)}{(2x+1)^2})dx=\int-\tfrac{-\tfrac{4x^3}{3}-2x^3-x-C_{1}}{(2x+1)^3} dx"


13 - step: Evaluate the integrals:


"\tfrac{(x+1)y(x)}{(2x+1)^2}=\tfrac{1}{3}(\tfrac{1}{4}(2x+1)+\tfrac{-6C_{1}+1}{8(2x+1)^2})+"C2

where C2 is an arbitrary constant.


14-step: Divide both sides by "\mu(x)=\tfrac{x+1}{(2x+1)^2}:"

"y(x)=\tfrac{(2x+1)^2(\tfrac{1}{12}(2x+1)+\tfrac{-6C_{1}+1}{24(2x+1)^2}+C_{2})}{x+1}"


15-step: Simplify the arbitrary constants:


"y(x)=\tfrac{(2x+1)^2(\tfrac{1}{12}(2x+1)+\tfrac{C_{1}}{(2x+1)^2}+C_{2})}{x+1}"


Answer:
"y(x) = \tfrac{C_{2}\sqrt{\smash[b]{2x+1}}x}{\sqrt{\smash[b]{-2x-1}}}+\tfrac{C_{1}\sqrt{\smash[b]{2x+1}}x}{\sqrt{\smash[b]{-2x-1}}(x+1)}+\tfrac{2x^3}{3(x+1)}+\tfrac{x^2}{2(x+1)}"